# FairJudge: MLLM Judging for Social-Attributes and Prompt–Image Alignment

### v1

@@ 1.  Introduction
Large-scale studies have documented representation disparities in T2I and vision–language models and have introduced diagnostic tests and scorecards for auditing them (Luccioni et al., 2023; Wan et al., 2024). Remedial approaches range from lightweight sampling control to small learned adaptors. Fair Diffusion steers the sampler toward user-specified demographic proportions without retraining the backbone (Friedrich et al., 2023); Fair Mapping inserts a linear adapter to systematically reduce demographic skew with minimal compute (Li et al., 2023). Generation-time prompting with chain-of-thought (CoT), as in FairCoT, leverages reasoning to mitigate bias during image synthesis (Al Sahili et al., 2024). While effective for mitigation, these approaches operate on the generator; they do not provide an evaluation protocol that is calibrated, evidence-grounded, and appropriately abstaining. On the evaluation side, contrastive similarity (e.g., CLIP) is a dominant proxy for text–image faithfulness (Radford et al., 2021; Hessel et al., 2021), and VQA-style checklists probe finer-grained grounding (Hu et al., 2023), but both can reward surface correlation and penalize uncertainty. Preference-trained rankers (ImageReward, PickScore) improve agreement with human choices (Xu et al., 2023; Kirstain et al., 2023) yet remain scalar scorers with limited transparency. Recent “LLM-as-a-judge” work reports high agreement with human raters but also reveals position and verbosity biases, and thus calls for careful protocol design and meta-evaluation (Zheng et al., 2023; Chen et al., 2024; Wang et al., 2025). In short: we lack a lightweight, evaluation-oriented procedure that (i) reasons explicitly about visual evidence, (ii) adheres to label constraints, (iii) abstains when cues are insufficient, and (iv) applies beyond face-visible demographics to attributes like religion, culture, disability, and profession.
@@ 2.1.  Bias in Text-to-Image Models
Text-to-image (T2I) systems inherit and sometimes amplify social biases present in their training data. Early large-scale analyses show systematic disparities across gender, skin tone, geo-culture, and professions, and propose diagnostic protocols and scorecards for auditing representation (Luccioni et al., 2023; Wan et al., 2024). Several debiasing methods have been explored post-hoc or with light adaptation: Fair Diffusion steers sampling toward user-specified demographic proportions (Friedrich et al., 2023), while Fair Mapping learns a small linear adaptor to systematically reduce demographic skew with minimal compute (Li et al., 2023). Concurrently, FairCoT leverages chain-of-thought (CoT) reasoning in multimodal LLMs to mitigate bias during generation-time prompting (Al Sahili et al., 2024). Our work differs in purpose: rather than intervening on the generator, we study evaluation fairness by treating multimodal LLMs (MLLMs) as judges over social attributes and alignment outcomes.
@@
Face-centric attribute benchmarks remain widely used to quantify gender, race, and age performance. FairFace provides a balanced dataset across seven race groups and age/gender labels (Kärkkäinen and Joo, 2021). For context-embedded protected-attribute cues beyond faces, the PaTA benchmark was curated alongside the DeAR debiasing method for VLMs (Seth et al., 2023). Tooling spans classic face-attribute models and modern vision-language recognition. In practice, many works rely on the open-source DeepFace toolkit to obtain age, gender, and race predictions (Serengil, 2024), and on CLIP (Radford et al., 2021) for zero-shot category attribution via text prompts. However, audits of CLIP-family encoders find systematic gender and racial skew whose magnitude depends on model size and, crucially, the pre-training data source/scale (Al Sahili et al., 2025). We extend this line by prompting MLLMs (GPT-4.1, Gemini, LLaMA-4) to directly infer social attributes, enabling coverage of non-facial categories (e.g., religion, culture, disability) with explanations.
@@
CLIP-based similarity remains a dominant automatic proxy for prompt faithfulness (Hessel et al., 2021). Beyond cosine-based scores, VQA-driven faithfulness frameworks such as TIFA generate question–answer checklists to probe fine-grained concept grounding (Hu et al., 2023). Preference-trained reward/score models (e.g., ImageReward and PickScore) leverage large-scale human comparisons to better match human judgments for T2I ranking and evaluation (Xu et al., 2023; Kirstain et al., 2023). Recently, X-IQE proposes a Chain-of-Thought, explanation-oriented evaluator using visual LLMs, reporting improved agreement with humans while outputting interpretable rationales (Chen et al., 2023). In our alignment study, we compare CLIP similarity against MLLM-based judging calibrated with an X-IQE–style scoring protocol to a standardized [1,1][-1,1] range.
@@ 2.3.  LLMs/MLLMs as Judges
Using LLMs as evaluators (“LLM-as-a-judge”) has become a practical alternative to costly human evaluation. Foundational work reports high agreement with human preferences but also identifies position, verbosity, and self-enhancement biases, motivating careful protocol design and meta-evaluation (Zheng et al., 2023). Subsequent studies investigate reliability, consistency, and bias in LLM judging, including on open-ended tasks (Chen et al., 2024). For images, recent work shows that MLLMs can serve as safety and policy judges with structured reasoning pipelines, revealing both promise and pitfalls for multimodal evaluation (Wang et al., 2025). Our work complements this literature by: (i) framing social-attribute prediction as a judging task on face and non-face domains (e.g., culture, religion, disability), (ii) quantifying per-attribute accuracy and generalization across FairFace, PaTA, and new human-centric scenarios, and (iii) benchmarking alignment judging on IdenProf professions and FairCoT—contrasting CLIP-based scores with MLLM judges that also supply free-form rationales.

### v2

@@ 1.  Introduction
Large-scale studies have documented representation disparities in T2I and vision–language models and have introduced diagnostic tests and scorecards for auditing them (luccioni2023stablebias; wan2024survey). Remedial approaches range from lightweight sampling control to small learned adaptors. Fair Diffusion steers the sampler toward user-specified demographic proportions without retraining the backbone (friedrich2023fairdiffusion); Fair Mapping inserts a linear adapter to systematically reduce demographic skew with minimal compute (li2023fairmapping). Generation-time prompting with chain-of-thought (CoT), as in FairCoT, leverages reasoning to mitigate bias during image synthesis (sahili2025faircot). While effective for mitigation, these approaches operate on the generator; they do not provide an evaluation protocol that is calibrated, evidence-grounded, and appropriately abstaining. On the evaluation side, contrastive similarity (e.g., CLIP) is a dominant proxy for text–image faithfulness (radford2021clip; hessel2021clipscore), and VQA-style checklists probe finer-grained grounding (hu2023tifa), but both can reward surface correlation and penalize uncertainty. Preference-trained rankers (ImageReward, PickScore) improve agreement with human choices (xu2023imagereward; kirstain2023pickapic) yet remain scalar scorers with limited transparency. Recent “LLM-as-a-judge” work reports high agreement with human raters but also reveals position and verbosity biases, and thus calls for careful protocol design and meta-evaluation (zheng2023llmasjudge; chen2024humansllms; wang2025mllmjudge). In short: we lack a lightweight, evaluation-oriented procedure that (i) reasons explicitly about visual evidence, (ii) adheres to label constraints, (iii) abstains when cues are insufficient, and (iv) applies beyond face-visible demographics to attributes like religion, culture, disability, and profession.
@@ 2.1.  Bias in Text-to-Image Models
Text-to-image (T2I) systems inherit and sometimes amplify social biases present in their training data. Early large-scale analyses show systematic disparities across gender, skin tone, geo-culture, and professions, and propose diagnostic protocols and scorecards for auditing representation (luccioni2023stablebias; wan2024survey). Several debiasing methods have been explored post-hoc or with light adaptation: Fair Diffusion steers sampling toward user-specified demographic proportions (friedrich2023fairdiffusion), while Fair Mapping learns a small linear adaptor to systematically reduce demographic skew with minimal compute (li2023fairmapping). Concurrently, FairCoT leverages chain-of-thought (CoT) reasoning in multimodal LLMs to mitigate bias during generation-time prompting (sahili2025faircot). Our work differs in purpose: rather than intervening on the generator, we study evaluation fairness by treating multimodal LLMs (MLLMs) as judges over social attributes and alignment outcomes.
@@
Face-centric attribute benchmarks remain widely used to quantify gender, race, and age performance. FairFace provides a balanced dataset across seven race groups and age/gender labels (karkkainen2021fairface). For context-embedded protected-attribute cues beyond faces, the PaTA benchmark was curated alongside the DeAR debiasing method for VLMs (seth2023dear). Tooling spans classic face-attribute models and modern vision-language recognition. In practice, many works rely on the open-source DeepFace toolkit to obtain age, gender, and race predictions (serengilDeepFace), and on CLIP (radford2021clip) for zero-shot category attribution via text prompts. However, audits of CLIP-family encoders find systematic gender and racial skew whose magnitude depends on model size and, crucially, the pre-training data source/scale (sahili2025data). We extend this line by prompting MLLMs (GPT-4.1, Gemini, LLaMA-4) to directly infer social attributes, enabling coverage of non-facial categories (e.g., religion, culture, disability) with explanations.
@@
CLIP-based similarity remains a dominant automatic proxy for prompt faithfulness (hessel2021clipscore). Beyond cosine-based scores, VQA-driven faithfulness frameworks such as TIFA generate question–answer checklists to probe fine-grained concept grounding (hu2023tifa). Preference-trained reward/score models (e.g., ImageReward and PickScore) leverage large-scale human comparisons to better match human judgments for T2I ranking and evaluation (xu2023imagereward; kirstain2023pickapic). Recently, X-IQE proposes a Chain-of-Thought, explanation-oriented evaluator using visual LLMs, reporting improved agreement with humans while outputting interpretable rationales (chen2023xiqe). In our alignment study, we compare CLIP similarity against MLLM-based judging calibrated with an X-IQE–style scoring protocol to a standardized [1,1][-1,1] range.
@@ 2.3.  LLMs/MLLMs as Judges
Using LLMs as evaluators (“LLM-as-a-judge”) has become a practical alternative to costly human evaluation. Foundational work reports high agreement with human preferences but also identifies position, verbosity, and self-enhancement biases, motivating careful protocol design and meta-evaluation (zheng2023llmasjudge). Subsequent studies investigate reliability, consistency, and bias in LLM judging, including on open-ended tasks (chen2024humansllms). For images, recent work shows that MLLMs can serve as safety and policy judges with structured reasoning pipelines, revealing both promise and pitfalls for multimodal evaluation (wang2025mllmjudge). Our work complements this literature by: (i) framing social-attribute prediction as a judging task on face and non-face domains (e.g., culture, religion, disability), (ii) quantifying per-attribute accuracy and generalization across FairFace, PaTA, and new human-centric scenarios, and (iii) benchmarking alignment judging on IdenProf professions and FairCoT—contrasting CLIP-based scores with MLLM judges that also supply free-form rationales.